I am having difficulty completing this question. Any help is appreciated. Thanks in advance.
Substitution Method:
x=2-y
y=3-2x

Answer by mananth(16949)

x=2-y
y=3-2x
y=3-2(2-y)
y=3-4+2y
y=-1+2y
-2y -2y
-y=-1
/-1
y=1
plug y in any equation to get x